OpenStudy (anonymous):

verify answer: write this as a single log : 2ln x - 4ln (1/y) - 3ln (xy) I get: ln (y/4x)

OpenStudy (blurbendy):

i think it's just ln(y/x)

Directrix (directrix):

@blurbendy How did you get ln(y/x) ? Thanks.

OpenStudy (anonymous):

\[2lnx - 4 \ln \frac{ 1 }{ y } - 3\ln (xy) \] \[lnx^2 -4 \left[ \ln1 - lny \right] -3\ln(xy)\] since ln1 =0 you get: \[lnx^2 +lny^4-lnx^3 -lny^3 \] \[\ln \frac{ x^2y^4 }{ x^3y^3}\] since you are dividing you subtract exponents in the x : 2- 3 = -1 therefore the x goes in the denominator and in the y : 4-3 =1 so you get \[\ln \frac{ y }{ x }\]

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

@camilasanchez Correct Answers without work are not helpful (my opinion).